Would someone please help me. I forgot how to do this kind of problem.
Use the foil method to find the product
(6x^2 - 4)(x^6 - 5)

Answer by stanbon(75887)

Use the foil method to find the product
(6x^2 - 4)(x^6 - 5)
----
FOIL:
F: (6x^2)(x^6) = 6x^8
O: (6x^2)(-5) = -30x^2
I: (-4)(x^6) = -4x^6
L: (-4)(-5) = 20
---
Combine the products to get:
6x^8 -4x^6 -30x^2 + 20
==============================
Cheers,
Stan H.